Hello! I'm trying to solve this problem using my BA II Plus Financial Calculator and I can't seem to figure out how to solve it by calculator.

The value of the following cash flows four years from today is $8,001.52. The interest rate is 4.8 percent. What is the value of the Year 3 cash flow?

Year Cash Flow 1 - $ 1,585

2 - 1,752

3 - ?

4 - 2,765

Multiple Choice Answers:

$1419.77

$1292.69

$1559.34

$1354.74

$1487.92
Answer #1

CF1 = 1,585

CF2 = 1,752

CF3 = Need to find

CF4 = 2,765

I = 4.8%

Value of cash flow after 4 years = 8,001.52

8,001.52 = 1,585 * (1.048)3 + 1,752 * (1.048)2 + CF3 * (1.048) + 2,765

8,001.52 = 1,824.37 + 1,924.23 + CF3 * (1.048) + 2,765

8,001.52 = 6,513.60 + CF3 * (1.048)

CF3 = (8,001.52 - 6,513.60)/ 1.048

CF3 = 1,419.77

Option A is correct